Remove grouping symbols and simplify
a(a-5b)-6a(a+b)=

Answer by ptaylor(2198)

a(a-5b)-6a(a+b)=
a^2-5ab-6a^2-6ab=
a^2-6a^2-5ab-6ab=
-5a^2-11ab

Hope this helps---ptaylor